# Detecting Where Effects Occur by Testing Hypotheses in Order

### v1

@@ 1 Where do block-level effects occur?
Consider a block-randomized experiment with 44 experimental blocks such as the Detroit Promise Program [undefq] where treatment was assigned to students of different academic cohorts by five community colleges. The overall test rejects the null of no effects. The policy-maker considering changes across all of Detroit’s community colleges now asks: where did the effects occur? Were they concentrated in a few sites, or spread evenly? The answer determines whether to replicate the intervention everywhere or investigate why it succeeded in some places and failed in others.
@@
Two conditions drive this result. First, a stopping rule: test a child hypothesis only after rejecting its parent. Second, valid tests: each individual test controls its false positive rate at level $\alpha$. We show that two conditions suffice for weak FWER control — control when all null hypotheses are true. For strong control, we analyze how power decay from data splitting limits error accumulation through the tree and, when needed, apply an adaptive $\alpha$-adjustment. This framework builds on the closed testing tradition of [undefn], [undefs] and the sequential structured testing of [undefe], [undefd], and [undefo], which we extend to randomization-based inference in block-randomized experiments.
@@
Two traditions of research address heterogeneous treatment effects. One estimates conditional average treatment effects as functions of observed covariates [undefaa, undeff, e.g.]. The other studies the distribution of individual effects directly, asking how many units benefited and by how much [undefh, undefl, e.g.]. Our work is closer to the second tradition: rather than asking which covariates moderate effects, we ask which experimental blocks or groups of blocks show detectable effects. Where covariate-based approaches focus on columns of the data matrix, we focus on rows — the sites, cohorts, and schools where effects did or did not occur. The approaches are complementary; detections could guide subsequent covariate-based investigation of why effects vary.
We focus on block-randomized experiments with two treatment arms and roughly continuous outcomes. The results generalize to binary outcomes and multi-arm trials directly, although we do not engage those designs in our simulations or application.
The paper proceeds as follows. Section 2 develops the bottom-up and top-down testing approaches, proves weak FWER control (Theorem 2), and illustrates the performance of the procedure using simulations across a wide range of tree sizes. Section 3 addresses strong FWER control by analyzing how power decay through the tree naturally limits error accumulation, and develops an adaptive $\alpha$-adjustment for settings where this natural gating is insufficient. We then demonstrate the method on data simulated to follow the Detroit Promise Program design (Section 3.2) and apply it to 25 block-randomized education trials fielded by the MDRC [undefa]. [^2] In the supplement we develop a test statistic sensitive to distributional differences beyond mean shifts, so that effects that are positive in some blocks and negative in others do not cancel at the root. This statistic draws on energy statistics [undefy] and multivariate permutation tests [undefw, undefk].
@@ 2.1 Testing in every block: the basic problem of multiple testing
Since each block contains multiple units assigned at random to treatment, an analyst could treat each block as a mini-experiment [undefc]. With 100 blocks, we could execute 100 tests of the null hypothesis of no treatment effects. We should also adjust the rejection criteria for those tests to protect the decision-makers from making too many false positive errors. Why adjust? Imagine that, in fact, there were no effects in any of the 100 blocks. If the tests are independent of one another, and we have set our rejection threshold at $\alpha=0.05$, then we know that the probability that at least one of the 100 unadjusted tests would yield a statistically significant result, $p<0.05$, in error is: $\mathbb{P}(\text{at least one }p<0.05)=(1-(1-0.05)^{100})\approx 0.99$. That is, if we did a test in each of the 100 blocks in an experiment in which no effects occurred in any block, we would virtually certainly reject this null in one or more blocks — and make an error when we interpret this rejection of "no effects" in those blocks as implying some true causal effect occurred in those blocks. This error rate of a collection of tests is often called the Family-wise Error Rate (FWER), and it can be controlled by an adjustment. For example, the Bonferroni adjustment suggests that if we refuse to reject the null hypothesis unless $\alpha_{\text{adjusted}}=p\leq 0.05/100=0.0005$, then we will have controlled the FWER because P(at least one p<0.05)=(1(10.05/100)100)0.05P($\text${at least one }p<0.05)=$(1-$$^{100})\approx 0.05$ [undefm]. [^3] Technically speaking, this explanation shows that dividing the rejection threshold by the number of tests, controls the FWER in the “weak sense”, which means it controls the FWER when there are no effects in any block. We will discuss strong and weak control of the FWER later in the paper.
@@ 2.2 Sequential Structured Testing and Weak Control of the FWER
Figure 1 shows that each experimental block is nested within a hierarchy. When hypotheses are nested in this way, they can be tested in a specific order under rules that avoid the severe power loss of bottom-up adjustments. This insight originates in the “closed testing” framework of [undefn], developed for structured experiments by [undefs, undefv]. We build especially on the general rules for sequential nested testing articulated by [undefe, undefd] and [undefo]. We explain the reasoning behind the structured testing approach here before turning to formal development, simulations, and application.
@@
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefe, undefd] and connects with frameworks for strictly nested testing that would focus on a single branch of a tree [undefs]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
@@
Each test at a node, if executed on its own, has a false positive rate no greater than $\alpha$. In block-randomized experiments, randomization-based tests satisfy this by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undeft, Chapter 2].
@@ 2.2.1 Simulation Study of Weak Control of the FWER to Illustrate the Proof
Given a specification of a complete $k$-ary tree using $k$ nodes per level and $L$ levels we drew $p$-values from uniform distributions. The valid tests condition (Condition 2) implies that the distribution of $p$-values across tests of the true null should be stochastically dominated by the uniform [undefm, Chapter 3]. That is, in many tests of a true null hypothesis, only 5% of $p$-values should be less than $p=0.05$, 10% should be less than $p=0.1$, etc….
@@ 3.1 Strong control via error load and adaptive adjustment
When the error load exceeds 1 — as it will with wide trees, shallow depth, or high root power — we develop an adaptive $\alpha$-adjustment in Appendix B.4 (Supplement). If the test statistic has an asymptotically Normal distribution — Appendix C (Supplement) introduces one that does, following [undefk] — then we can compute power at each depth before testing begins and adjust $\alpha$ accordingly (Remark 8 and Theorem 9) as follows:
@@ 3.2 An Example: A Simulated Version of The Detroit Promise Program
The Detroit Promise Program (DPP) [undefq] involved enhanced advising, financial support, and encouragement of full-time enrollment to community college students in an effort to increase graduation rates. The trial randomized this bundle of treatments within five community colleges in the Detroit area: Henry Ford Community College, Macomb Community College, Oakland Community College, Schoolcraft College, and Wayne County Community College District. Each college fielded the trial for 3 academic year cohorts, and within the cohorts students were further subdivided into 1 to 4 experimental blocks within which the intervention was randomly assigned. To illustrate the algorithm, we generated a design that preserved the DPP structure — 5 college nodes under the overall node with varying numbers of cohorts and sub-cohort blocks, for 44 blocks total — but differed in ways that facilitated simulation. Every block had 50 students, and potential outcomes under control were drawn from $N(10,9)$ to roughly match the observed distribution of credits taken, but with more variation. We concentrated all non-zero effects in a single college: 9 blocks within HFCC have treatment effects while the 35 blocks across the other four colleges are pure null. This concentration mimics a pattern common in multi-site trials and visible in the MDRC data (Section 4): effects that are localized within particular sites rather than uniformly spread. Each block contains 50 students. For the illustration we use a large effect (Cohen’s $d=0.80$) within each of the 9 blocks with non-null effects so that the root-level test rejects and the tree structure is visible; the simulation study that follows evaluates realistic effect sizes.
@@
Figure 5: Results of top-down testing in a simulation of 44 experimental blocks following the pre-specified experimental design of the Detroit Promise Program [undefq]. Nine blocks within HFCC have non-zero effects (Cohen’s $d=0.80$); all other blocks are pure null. The algorithm identifies HFCC and descends into its cohorts and blocks while pruning null colleges. Abbreviations: Henry Ford Community College (HFCC), Macomb Community College (MCC), Oakland Community College (OCC), Schoolcraft College (SC), Wayne County Community College District (WCC). Blue nodes have non-zero causal effects.
This illustration shows that a policy maker seeking to identify which colleges or blocks showed treatment effects would find the bottom-up approach disappointing: the [undefj] adjustment across 44 blocks detects just one block despite nine having genuine effects. The structured approach leads attention away from three of the five colleges and identifies not only more individual blocks but also higher-level groupings — the HFCC college itself — that can direct further exploration toward understanding why effects concentrated there.
@@ 4 Application: The MDRC RCT Data
From 2003 to 2019 the MDRC organization worked with community colleges to field randomized field trials assessing the effects of interventions like financial support or improved advising meant to improve the educational outcomes of community college students. MDRC deposited data from 31 of these studies in the ICPSR data repository, where they are available under restricted data use agreements [undefa]. We applied the method to the 25 studies that randomly assigned an intervention to individual students within strata, following each study’s pre-specified series of splits. The outcome in each study is the number of community college credits taken in the first main session after the intervention — for example, credits taken during the next full academic year [undefab, undefu].
@@ 5 Limitations
The structured testing approach requires a test at each node that can detect any departure from the null, not just a mean shift. If half of the blocks have a large negative effect and half a large positive effect, a difference-in-means test at the root could fail to reject. [^8] Thanks to Kevin Quinn for this question. We address this by creating a test statistic based on energy statistics [undefr, undefx] combined with the multivariate permutation framework of [undefk] and related to the $d^{2}$-test of [undefg]. This statistic is sensitive to distributional differences of any kind and is described in Appendix C (Supplement) and implemented in the manytestsr R package. [^9] The manytestsr package is available at https://bowers-illinois-edu.github.io/manytestsr/. The combined Stephenson rank statistic of [undefl] is a promising alternative that merits investigation. [^10] The R package for that statistic is at https://bowers-illinois-edu.github.io/CMRSS.
@@
This paper focuses on FWER control. An alternative is to allocate the error budget dynamically as tests descend the tree, targeting control of the false discovery rate (FDR) rather than the FWER. The alpha-investing procedure of [undefb] and its refinements [undefp, undefz] provide a natural framework for this extension. FDR-oriented tree testing would offer different guarantees and different power profiles and deserves its own treatment.
@@ A.1 Conditions on the Procedure
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefad, undefd]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
@@
In block-randomized experiments, randomization-based tests satisfy the Valid Tests Rule by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undeft, § 2.4]. All tests that we consider in this paper are randomization-based tests justified by the design of the randomized experiments to which they are applied.
@@ A.2 Proof: Weak Control of the FWER Using Two Conditions
[undefs] establishes the analogous result for linear sequences of hypotheses: testing in a fixed order and stopping at the first non-rejection controls the FWER at $\alpha$. His argument is essentially the same gating logic, but applies only to chains, not trees with branching.
[undefo] proves FWER control for tree-structured hypotheses via a different route entirely: a Bonferroni bound over disjoint maximal true-null clusters. That proof exploits the hierarchy to partition the null hypotheses, rather than conditioning on the root.
@@ B.4 Adaptive Alpha Adjustment When Gating Is Insufficient
This approach requires, in [undefs]’s words, a “first true null hypothesis”, or an ordering of hypotheses such that non-rejection of the first hypothesis ought to imply a non-rejection of the subsequent hypotheses. This requirement raises questions about the test statistic used. A difference-of-means test statistic has power to detect shifts in the mean. But such statistics lose power in the presence of outliers [undefac], so a test on a smaller subset without outliers could have more power than a test on the full sample. Moreover, mean-difference statistics record the sign of effects: if half the blocks have positive effects and half negative, the overall mean difference may be near zero despite genuine effects in every block. In either case, the root-level test could fail to reject even when block-level effects are real, preventing the procedure from descending into the tree.
@@ C.1 A Combined Energy Statistic
We introduce a test statistic that captures not only mean differences but also distributional differences between treated and control units, in ways that are not fooled by strong countervailing effects. The strategy is to transform each unit’s outcome into several representations — ranks, pairwise distances [undefx], and nonlinear transforms — and then combine them into a single test. Specifically, we compute the following six scores for each unit:
@@
    The mean Euclidean distance between a unit’s outcome and the outcomes of all other units in the block, following the energy-distance framework of [undefy, undefx, undefr].
@@
    The hyperbolic tangent (tanh) transform of the outcome, which serves as an alternative to the log transform when the variable has many zeros [undefaf].
Each unit thus receives six scores, each capturing a different aspect of the outcome distribution. We combine these scores into a single test statistic and compare the observed value to its randomization distribution, using either large-sample chi-square approximations [undefg, undefae, undefw] or exact permutation tests when the sample size is small.
Why should this somewhat arbitrary combined test statistic have large-sample approximations? Under the null hypothesis of no effects and given the known randomization (within blocks in this case), a linear combination of mean differences follows a chi-square distribution. The required variances and covariances are functions of the randomization distribution, which the null and the design determine [undefg, undefw]. The independence_test function in the coin R package [undefk] implements this combination.

### v2

# arXiv:2602.21068v2[stat.ME] 27 Feb 2026

@@ 1 Where do block-level effects occur?
Consider a block-randomized experiment with 44 experimental blocks such as the Detroit Promise Program [undefr] where treatment was assigned to students of different academic cohorts by five community colleges. The overall test rejects the null of no effects. The policy-maker considering changes across all of Detroit’s community colleges now asks: where did the effects occur? Were they concentrated in a few sites, or spread evenly? The answer determines whether to replicate the intervention everywhere or investigate why it succeeded in some places and failed in others.
@@
Two conditions drive this result. First, a stopping rule: test a child hypothesis only after rejecting its parent. Second, valid tests: each individual test controls its false positive rate at level $\alpha$. We show that two conditions suffice for weak FWER control — control when all null hypotheses are true. For strong control, we analyze how power decay from data splitting limits error accumulation through the tree and, when needed, apply an adaptive $\alpha$-adjustment. This framework builds on the closed testing tradition of [undefo], [undeft] and the sequential structured testing of [undefe], [undefd], and [undefp], which we extend to randomization-based inference in block-randomized experiments.
@@
Two traditions of research address heterogeneous treatment effects. One estimates conditional average treatment effects as functions of observed covariates [undefab, undeff, e.g.]. The other studies the distribution of individual effects directly, asking how many units benefited and by how much [undefh, undefm, e.g.]. Our work is closer to the second tradition: rather than asking which covariates moderate effects, we ask which experimental blocks or groups of blocks show detectable effects. Where covariate-based approaches focus on columns of the data matrix, we focus on rows — the sites, cohorts, and schools where effects did or did not occur. The approaches are complementary; detections could guide subsequent covariate-based investigation of why effects vary. A related question arises when sites differ not only in treatment effects but in local ecological conditions that confound between-site comparisons. [undefk] develop methods to isolate organizational effectiveness from such confounding in multisite trials using parametric models and empirical bayes style statistical inference.
We focus on block-randomized experiments with two treatment arms and roughly continuous outcomes using randomization-based statistical inference. The results generalize to binary outcomes and multi-arm trials directly, although we do not engage those designs in our simulations or application.
The paper proceeds as follows. Section 2 develops the bottom-up and top-down testing approaches, proves weak FWER control (Theorem 2), and illustrates the performance of the procedure using simulations across a wide range of tree sizes. Section 3 addresses strong FWER control by analyzing how power decay through the tree naturally limits error accumulation, and develops an adaptive $\alpha$-adjustment for settings where this natural gating is insufficient. We then demonstrate the method on data simulated to follow the Detroit Promise Program design (Section 3.2) and apply it to 25 block-randomized education trials fielded by the MDRC [undefa]. [^2] In the supplement we develop a test statistic sensitive to distributional differences beyond mean shifts, so that effects that are positive in some blocks and negative in others do not cancel at the root. This statistic draws on energy statistics [undefz] and multivariate permutation tests [undefx, undefl].
@@ 2.1 Testing in every block: the basic problem of multiple testing
Since each block contains multiple units assigned at random to treatment, an analyst could treat each block as a mini-experiment [undefc]. With 100 blocks, we could execute 100 tests of the null hypothesis of no treatment effects. We should also adjust the rejection criteria for those tests to protect the decision-makers from making too many false positive errors. Why adjust? Imagine that, in fact, there were no effects in any of the 100 blocks. If the tests are independent of one another, and we have set our rejection threshold at $\alpha=0.05$, then we know that the probability that at least one of the 100 unadjusted tests would yield a statistically significant result, $p<0.05$, in error is: $\mathbb{P}(\text{at least one }p<0.05)=(1-(1-0.05)^{100})\approx 0.99$. That is, if we did a test in each of the 100 blocks in an experiment in which no effects occurred in any block, we would virtually certainly reject this null in one or more blocks — and make an error when we interpret this rejection of "no effects" in those blocks as implying some true causal effect occurred in those blocks. This error rate of a collection of tests is often called the Family-wise Error Rate (FWER), and it can be controlled by an adjustment. For example, the Bonferroni adjustment suggests that if we refuse to reject the null hypothesis unless $\alpha_{\text{adjusted}}=p\leq 0.05/100=0.0005$, then we will have controlled the FWER because P(at least one p<0.05)=(1(10.05/100)100)0.05P($\text${at least one }p<0.05)=$(1-$$^{100})\approx 0.05$ [undefn]. [^3] Technically speaking, this explanation shows that dividing the rejection threshold by the number of tests, controls the FWER in the “weak sense”, which means it controls the FWER when there are no effects in any block. We will discuss strong and weak control of the FWER later in the paper.
@@ 2.2 Sequential Structured Testing and Weak Control of the FWER
Figure 1 shows that each experimental block is nested within a hierarchy. When hypotheses are nested in this way, they can be tested in a specific order under rules that avoid the severe power loss of bottom-up adjustments. This insight originates in the “closed testing” framework of [undefo], developed for structured experiments by [undeft, undefw]. We build especially on the general rules for sequential nested testing articulated by [undefe, undefd] and [undefp]. We explain the reasoning behind the structured testing approach here before turning to formal development, simulations, and application.
@@
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefe, undefd] and connects with frameworks for strictly nested testing that would focus on a single branch of a tree [undeft]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
@@
Each test at a node, if executed on its own, has a false positive rate no greater than $\alpha$. In block-randomized experiments, randomization-based tests satisfy this by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undefu, Chapter 2].
@@ 2.2.1 Simulation Study of Weak Control of the FWER to Illustrate the Proof
Given a specification of a complete $k$-ary tree using $k$ nodes per level and $L$ levels we drew $p$-values from uniform distributions. The valid tests condition (Condition 2) implies that the distribution of $p$-values across tests of the true null should be stochastically dominated by the uniform [undefn, Chapter 3]. That is, in many tests of a true null hypothesis, only 5% of $p$-values should be less than $p=0.05$, 10% should be less than $p=0.1$, etc….
@@ 3.1 Strong control via error load and adaptive adjustment
When the error load exceeds 1 — as it will with wide trees, shallow depth, or high root power — we develop an adaptive $\alpha$-adjustment in Appendix B.4 (Supplement). If the test statistic has an asymptotically Normal distribution — Appendix C (Supplement) introduces one that does, following [undefl] — then we can compute power at each depth before testing begins and adjust $\alpha$ accordingly (Remark 8 and Theorem 9) as follows:
@@ 3.2 An Example: A Simulated Version of The Detroit Promise Program
The Detroit Promise Program (DPP) [undefr] involved enhanced advising, financial support, and encouragement of full-time enrollment to community college students in an effort to increase graduation rates. The trial randomized this bundle of treatments within five community colleges in the Detroit area: Henry Ford Community College, Macomb Community College, Oakland Community College, Schoolcraft College, and Wayne County Community College District. Each college fielded the trial for 3 academic year cohorts, and within the cohorts students were further subdivided into 1 to 4 experimental blocks within which the intervention was randomly assigned. To illustrate the algorithm, we generated a design that preserved the DPP structure — 5 college nodes under the overall node with varying numbers of cohorts and sub-cohort blocks, for 44 blocks total — but differed in ways that facilitated simulation. Every block had 50 students, and potential outcomes under control were drawn from $N(10,9)$ to roughly match the observed distribution of credits taken, but with more variation. We concentrated all non-zero effects in a single college: 9 blocks within HFCC have treatment effects while the 35 blocks across the other four colleges are pure null. This concentration mimics a pattern common in multi-site trials and visible in the MDRC data (Section 4): effects that are localized within particular sites rather than uniformly spread. Each block contains 50 students. For the illustration we use a large effect (Cohen’s $d=0.80$) within each of the 9 blocks with non-null effects so that the root-level test rejects and the tree structure is visible; the simulation study that follows evaluates realistic effect sizes.
@@
Figure 5: Results of top-down testing in a simulation of 44 experimental blocks following the pre-specified experimental design of the Detroit Promise Program [undefr]. Nine blocks within HFCC have non-zero effects (Cohen’s $d=0.80$); all other blocks are pure null. The algorithm identifies HFCC and descends into its cohorts and blocks while pruning null colleges. Abbreviations: Henry Ford Community College (HFCC), Macomb Community College (MCC), Oakland Community College (OCC), Schoolcraft College (SC), Wayne County Community College District (WCC). Blue nodes have non-zero causal effects.
This illustration shows that a policy maker seeking to identify which colleges or blocks showed treatment effects would find the bottom-up approach disappointing: the [undefj] adjustment across 44 blocks detects just one block despite nine having genuine effects. The structured approach leads attention away from three of the five colleges and identifies not only more individual blocks but also higher-level groupings — the HFCC college itself — that can direct further exploration toward understanding why effects concentrated there [undefk, see].
@@ 4 Application: The MDRC RCT Data
From 2003 to 2019 the MDRC organization worked with community colleges to field randomized field trials assessing the effects of interventions like financial support or improved advising meant to improve the educational outcomes of community college students. MDRC deposited data from 31 of these studies in the ICPSR data repository, where they are available under restricted data use agreements [undefa]. We applied the method to the 25 studies that randomly assigned an intervention to individual students within strata, following each study’s pre-specified series of splits. The outcome in each study is the number of community college credits taken in the first main session after the intervention — for example, credits taken during the next full academic year [undefac, undefv].
@@ 5 Limitations
The structured testing approach requires a test at each node that can detect any departure from the null, not just a mean shift. If half of the blocks have a large negative effect and half a large positive effect, a difference-in-means test at the root could fail to reject. [^8] Thanks to Kevin Quinn for this question. We address this by creating a test statistic based on energy statistics [undefs, undefy] combined with the multivariate permutation framework of [undefl] and related to the $d^{2}$-test of [undefg]. This statistic is sensitive to distributional differences of any kind and is described in Appendix C (Supplement) and implemented in the manytestsr R package. [^9] The manytestsr package is available at https://bowers-illinois-edu.github.io/manytestsr/. The combined Stephenson rank statistic of [undefm] is a promising alternative that merits investigation. [^10] The R package for that statistic is at https://bowers-illinois-edu.github.io/CMRSS.
@@
This paper focuses on FWER control. An alternative is to allocate the error budget dynamically as tests descend the tree, targeting control of the false discovery rate (FDR) rather than the FWER. The alpha-investing procedure of [undefb] and its refinements [undefq, undefaa] provide a natural framework for this extension. FDR-oriented tree testing would offer different guarantees and different power profiles and deserves its own treatment.
@@ A.1 Conditions on the Procedure
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefae, undefd]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
@@
In block-randomized experiments, randomization-based tests satisfy the Valid Tests Rule by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undefu, § 2.4]. All tests that we consider in this paper are randomization-based tests justified by the design of the randomized experiments to which they are applied.
@@ A.2 Proof: Weak Control of the FWER Using Two Conditions
[undeft] establishes the analogous result for linear sequences of hypotheses: testing in a fixed order and stopping at the first non-rejection controls the FWER at $\alpha$. His argument is essentially the same gating logic, but applies only to chains, not trees with branching.
[undefp] proves FWER control for tree-structured hypotheses via a different route entirely: a Bonferroni bound over disjoint maximal true-null clusters. That proof exploits the hierarchy to partition the null hypotheses, rather than conditioning on the root.
@@ B.4 Adaptive Alpha Adjustment When Gating Is Insufficient
This approach requires, in [undeft]’s words, a “first true null hypothesis”, or an ordering of hypotheses such that non-rejection of the first hypothesis ought to imply a non-rejection of the subsequent hypotheses. This requirement raises questions about the test statistic used. A difference-of-means test statistic has power to detect shifts in the mean. But such statistics lose power in the presence of outliers [undefad], so a test on a smaller subset without outliers could have more power than a test on the full sample. Moreover, mean-difference statistics record the sign of effects: if half the blocks have positive effects and half negative, the overall mean difference may be near zero despite genuine effects in every block. In either case, the root-level test could fail to reject even when block-level effects are real, preventing the procedure from descending into the tree.
@@ C.1 A Combined Energy Statistic
We introduce a test statistic that captures not only mean differences but also distributional differences between treated and control units, in ways that are not fooled by strong countervailing effects. The strategy is to transform each unit’s outcome into several representations — ranks, pairwise distances [undefy], and nonlinear transforms — and then combine them into a single test. Specifically, we compute the following six scores for each unit:
@@
    The mean Euclidean distance between a unit’s outcome and the outcomes of all other units in the block, following the energy-distance framework of [undefz, undefy, undefs].
@@
    The hyperbolic tangent (tanh) transform of the outcome, which serves as an alternative to the log transform when the variable has many zeros [undefag].
Each unit thus receives six scores, each capturing a different aspect of the outcome distribution. We combine these scores into a single test statistic and compare the observed value to its randomization distribution, using either large-sample chi-square approximations [undefg, undefaf, undefx] or exact permutation tests when the sample size is small.
Why should this somewhat arbitrary combined test statistic have large-sample approximations? Under the null hypothesis of no effects and given the known randomization (within blocks in this case), a linear combination of mean differences follows a chi-square distribution. The required variances and covariances are functions of the randomization distribution, which the null and the design determine [undefg, undefx]. The independence_test function in the coin R package [undefl] implements this combination.